An open cone has a slant height of 5 cm and a curved surface area of 251cm square .find the radius of its circular base leaving your answer corrects to 3 significant figures

Answer by addingup(3677)

251 = Pi(r)5
251 = 15.71(r)
r = 251/15.71 = 15.977